Skip to main content

Basic Concepts of Probability and Binomial Distribution , Poisson Distribution.

 Probability: 

Basic concepts of Probability: 

Probability is a way to measure hoe likely something is to happen. Probability is number between 0 and 1, where probability is 0 means is not happen at all and probability is 1 means it will be definitely happen, e.g. if we tossed coin there is a 50% chance to get head and 50% chance to get tail, it can be represented in probability as 0.5 for each outcome to get head and tail. Probability is used to help us taking decision and predicting the likelihood of the event in many areas, that are science, finance and Statistics. 

Now we learn the some basic concepts that used in Probability: 

i) Random Experiment OR Trail: A Random Experiment is an process that get one or more possible outcomes. examples of random experiment include tossing a coin, rolling a die, drawing  a card from pack of card etc. using this we specify the possible outcomes known as sample pace. 

ii)Outcome: An outcome is a result of experiment. an outcome is one of the possible result that can occur when random experiment is performed. e.g. if we tossed coin the possible outcomes are either head or tail, that outcome of experiment of tossing coin (i.e. H, T). the some of probability of all possible outcome of an experiment is always equal to 1.

iii) Sample Space: Sample Space is the set of all possible outcome that occur in experiment. it is denoted  as "S". e.g. if we tossed coin the sample space consist the following possible outcomes as head and tail. (i.e. H, T), in experiment of rolling die following are possible outcomes as 1,2,3,4,5,6.

iv) Event : event means we are interested in getting particular type of outcome out off set of all type of outcomes. e.g. in coin tossing experiment we get H & T but we are interested to getting H, it is getting H is event. 

v)Certain Event: A certain event is an event that guaranteed to occur. or if an event  A is the entire sample space then the event A is called certain event. the probability of certain event is 1. 

vi)Mutually Exclusive: Mutually Exclusive events are those event that cannot occur at the same time. if two events are said to be mutually exclusive event means one event can occur then other is not occur. or  if two events are said to be mutually exclusive event if both event cannot happen simultaneously , e.g. tossing a coin we get only head or tail only at one time both not happen same time.

vii)Equally Likely Event: Equally likely event are event that have same probability of occurrence. e.g. the probability of rolling a die we get 1,2,3,4,5,6 and each have same chance out occur or each have same probability.

viii) Impossible Event: the event dose not have any sample point is called impossible event. or there is no chance to occur this event. e.g. we rolling die then it is impossible to get 7. it is an impossible event. probability of impossible event is always 0.

ix)Exhaustive Event: A event is said to be exhaustive event if an event include all possible outcomes of an experiment. it covers all possible outcomes of experiment. e.g. if we toss coin two exhaustive events are Head and Tail. there is no other outcome of that experiment. 

x) Union of Two Event:  The event containing the sample point which are in either in event A or in event B or in both is called the union of the event A and B. And it is denoted as AUB Or A+B .

Example: if A = { 1, 2, 4 } and B = { 3, 5 }

then AUB = { 1, 2, 3, 4, 5 }

Note that Union of the event A and B means Combining the  all point in the both event A and B.  

xi ) Intersection of Two event:  The event containing sample point of event A and event B both is called Intersection of the event A and B. and it is denotes as  A⋂B.

Example: if A = { 2, 4, 8 } and B= { 2, 8, 6 , 10 }

therefore A⋂B = { 2, 8 }

Note: Intersection of  Event A and B means collecting common point in the both events A and B. 

 Definition and Properties of Probability:

Definition: if S is a sample space and it has n sample points they are equally likely, mutually exclusive and exhaustive and an event A has m points then the ration m/n is called Probability of event A and it is denoted as P(A). 

i.e. 

P(A) = (m/n) = (Number of point in event A / Number of point in sample space S).


Properties: 1. Probability of certain event is 1. i.e. P(S) = 1

(Certain Event: A certain event is an event that guaranteed to occur. or if an event  A is the entire sample space then the event A is called certain event. the probability of certain event is 1. )

Proof: If S has n points  and A is certain event then it contain entire sample space S or sample Point, then A has n point. therefore by definition of probability,

P(A) = (Number of point in A / Number of point in S) = (n/n) = 1


2.Probability od Impossible event is zero. i.e. P(ф) = 0 (Impossible event is denoted as ф)

Proof: we known Impossible event has no point. 

P(ф) = (0/n) = 0

(Impossible Event: the event dose not have any sample point is called impossible event. or there is no chance to occur this event. e.g. we rolling die then it is impossible to get 7. it is an impossible event. probability of impossible event is always 0.)


3. Probability of an event always lies between 0 and 1. i.e. 0≤ P(A) ≥ 1

Proof: if S has n point and A has m points then the number of points in event A is less than or equal to n because S is sample space it contains all point of events. and point in A is greater than 0, 

0≤ m≤ n

now dividing by n, we get

0≤  (m/n) ≤  (n/n)

0≤  P(A) ≤ 1.

Hence, Probability of an event always lies between 0 and 1

Example 1. : find the probability of getting tail in coin tossing event.

Answer: In a coin tossing experiment there are only two outcomes that are head and tail, therefore the sample space is two pints head and tail only.  consider event A is getting tail. therefore the event A has only one point. i.e. n(S)= n = 2 and n(A)= m=1. 

    P(A) = (m /n) = (1/2)= 0.5 

Therefore the probability of getting tail is Equal to  0.5


Example 2. An urn contains 8 white balls and 12 black balls. find the probability of drawing a white ball.

Answer: 

we know the definition of probability of an event A is  P(A) = (m /n)

where s is all points in sample space, here 8 white and 12 black balls therefore the total balls are           8 + 12  = 20   i.e. total balls are 20 means sample space has 20 balls. n = 20

and m is the point in event here getting a white ball is an event A therefore the 8 white balls means       n = 8  

the Probability of drawing a white ball = P ( A ) = ( m / n )  = 8 / 20  = 0.4

Hence the probability of drawing a white ball is  equal to 0.4 


Example 3. Find the Probability that a card drawn from pack of cards is King.

Answer: 

here the event A is getting king card 

 We know the definition of probability of an event A is  P(A) = (m /n)

the total cards in pack of card is 52 i.e. n = 52

and total king card in pack of a card is 4 i.e. m = 4

the probability of getting king card = P(A) = (m /n)

P(A) = (m /n) = (4 / 52 ) = 1 / 13 = 0.0769

The Probability of drawing a card from pack of card is king is equal to   0.0769 


Example 4.  Find the Probability of getting 3 or 5 in a throwing a dice.

Answer: the event  A is that getting 3 or 5 in throwing dice.

therefore the sample space of throwing a dice is s = { 1, 2, 3, 4, 5, 6}

the total points in sample space is 6 i.e. n= 6

and the getting 3 or 5 means m = 2 

the probability od getting 3 or 5 is = P(A) = (m /n)  =  (2 / 6 )

P(A) = ( 1 / 3 ) = 0.3333

The Probability of getting 3 or 5 in throwing a dice is equal to  0.3333

 Probability Examples:

Example 1. An urn contains 10 white and 5 black ball's. two ball are drawn from urn without replacement. find the probability that i) both ball will be of same colour. ii) both ball will be of different colours.

Answer : Given an urn contains 10 white & 5 black balls, total number of ball's is 15

i) Probability that both ball's will be of same colour.

here given that two colours of balls, therefore we consider firstly the probability of drawing two ball's is white balls without replacement. 

the probability of drawing a white ball in first drawn is 10/15. (i.e. if we want to find probability of drawing white ball is calculated as number of white ball's divided by total number of ball's = 10/15). in this case we- select next ball without replacement therefore first ball is not replaced then 9 white ball's are left in urn and total number of ball's is 14. the probability of drawing second ball is white is 9/14. (i.e. if we want to find probability of drawing white ball is calculated as number of white ball's divided by total number of ball's = 9/14). 

hence the probability of drawing both ball's is white without replacement is = (10/15) x (9/14) =3/7.

now we consider the drawing both ball's is black ball's.

he probability of drawing a blackball in first drawn is 5/15. (i.e. if we want to find probability of drawing black ball is calculated as number of black ball's divided by total number of ball's = 5/15). in this case we select next ball without replacement therefore first ball is not replaced then 4 black ball's are left in urn and total number of ball's is 14. the probability of drawing second ball is black is 4/14. (i.e. if we want to find probability of drawing white ball is calculated as number of white ball's divided by total number of ball's = 4/14). 

hence the probability of drawing both ball's is white without replacement is = (5/15) x (4/14) =1/21.

therefore The probability that drawing two ball's is same colour without replacement is sum of both ball's is white and black ball's is = (3/7) + (1/21) = 10/21.

ii) probability of drawing two ball's will be different in colours. 

for finding this probability the probability of white ball drawing in first drawn is 10/15

and the probability of black ball drawing in second drawn is 5/14.

therefore probability of  drawing white ball in first drawn and black ball in second drawn                      is = (10/15) x (5/14) = 0.23

similarly the probability of black ball drawing in first drawn is 5/15

and the probability of white ball drawing in second drawn is 10/14.

therefore probability of  drawing black ball in first drawn and white ball in second drawn                      is = (5/15) x (10/14) = 0.23

therefore the probability of drawing both ball's have different colour without replacement is equal to sum of  probability of  drawing black ball in first drawn and white ball in second drawn  and probability of  drawing white ball in first drawn and black ball in second drawn is = 0.23+0.23 =0.46

Therefore the probability of drawing both ball's have different colour without replacement is = 0.46


Example 2. A Box contains 20 tickets numbered as 1 to 20. a ticket drawn randomly  from the box. find the probability of ticket will be: i) multiple of 3, ii) a multiple of 5, iii) a multiple of 3 or 5,           iv) a ticket multiple of 3 and five.

Answer:  Given that a box contains 20 tickets and is numbered as 1 to 20,

i) Probability that the numbers on ticket will be a multiple of 3.

the ticket have 1 to 20 number in which the number on ticket will be multiple of 3 is 3,6,9,12,15,18.

number on tickets which are multiple of 3 is 6 numbers therefore the probability that the number on ticket will be a multiple of 3 is = 6/20

(i.e. 6 numbers are multiple of 3  and total number of tickets is 20).

Probability that the numbers on ticket will be a multiple of 3 is 6/20

ii) Probability that the number on ticket will be a multiple of 5.

the ticket have 1 to 20 number in which the numbers on ticket will be multiple of 5 is 5,10,15,20

number on tickets which are multiple of 5 is 4 numbers therefore the probability that the number on ticket will be a multiple of 5 is = 4/20

(i.e. 4 numbers are multiple of 5  and total number of tickets is 20)

Probability that the numbers on ticket will be a multiple of 5 is 4/20

iii)  Probability that the number on ticket will be a multiple of 3 or 5.

the ticket have 1 to 20 number in which the numbers on ticket will be multiple of 3 or 5 is 3,6,9,12,15,18,5,10,20.

number on tickets which are multiple of 3  or 5 is 9 numbers therefore the probability that the number on ticket will be a multiple of 3 or 5  is = 9/20

(i.e. 4 numbers are multiple of 5  and 6 numbers are multiple of 3  15 is common therefore 9 numbers are multiple of 3 or 5, and total number of tickets is 20)

Probability that the numbers on ticket will be a multiple of 3 or 5 is 9/20.

iv)  Probability that the number on ticket will be a multiple of 3 and 5.

the ticket have 1 to 20 number in which the numbers on ticket will be multiple of 3 and 5 is 15.

number on tickets which are multiple of 3  and 5 is 1 numbers therefore the probability that the number on ticket will be a multiple of 3 and 5  is = 1/20

(i.e. 4 numbers are multiple of 5  and 6 numbers are multiple of 3,  15 is common therefore 1 numbers are multiple of 3 and  5, and total number of tickets is 20)

Probability that the numbers on ticket will be a multiple of 3 and 5 is 1/20.

. 

 Probability Distribution: Binomial Distribution

In this article we learn binomial distribution, but for that we see some terms:

Probability Distribution: The way in which Probabilities are distributed over the variable corresponding to values of x is called Probability Distribution.

if  x is discrete random variable with probabilities Pi then the probability distribution is called discrete distribution function. if P(Xi ) is satisfying the condition that 0 ≤ P ≤ 1for all values od x and the             Σ Pi = 1 then P(X) is called probability density function.

Permutation: A group of some or all given number of things are arranged with considering their order is called permutation. and it is denoted as nPr

it is formulated as nPr   =n! /(n-r)! 

where ! - is call factorial ,

n! = n x (n-1)x (n-2)x .........x1

e.g. 5! = 5x4x3x2x1 =120

 Combination: A group of some or all given number of things are arranged without considering their order is called permutation. and it is denoted as  nCr

it is formulated as nCr  = (n!)/(n-r)! x r!

note that nCn  =1,  nC0  =1

e.g. the number of ways in which 11 students are selected from 16. is 

 16C11  = (16!)/(16-11)! x 11!

        =   (16x15x14x13x12x11!)/(5! x 11!)

        = (16x15x14x13x12) / (5x4x3x2x1)

        = 4368.


Binomial Distribution:

Definition: If an experiment which has only two results they can be expressed as Success and Failure &  such that p+q= 1 and this experiment is repeated n times then Probability of x success is P(X = x).

 P(X = x) = nCx  px qn-x  

here x is number of success then it takes values from 0 to n as x= 0,1,2,........n. and it is expressed as

P(X= x) = nCx  px qn-x       x = 0,1,2,..........n.

              =  0             otherwise

the resulting probabilities corresponding to values of x is called Binomial Distribution. 

Properties Of Binomial Distribution: 

i) Mean of binomial distribution is np

ii) Variance of binomial distribution id npq

iii) When we use binomial distribution: the trials are repeated n times. and n is finite. each trial has only two results. sum of Probability of  Success is p  and  Probability of Failure is q is always 1. 


Examples on Binomial Distribution.

Example 1. Define Binomial Distribution, State it Mean and Variance, If a random variable X has Binomial Distribution with parameter n = 4 and p = 2/3 = 0.6666.

find  i. Mean And Variance     ii. P(X=2),     iii. P(X<4).

Solution: The definition of Binomial Distribution is  If an experiment which has only two results they can be expressed as Success and Failure &  such that p+q= 1 and this experiment is repeated n times then Probability of x success is P(X = x).

 P(X = x) = nCx  px qn-x  

here x is number of success then it takes values from 0 to n as x= 0,1,2,........n. and it is expressed as

P(X= x) = nCx  px qn-x       x = 0,1,2,..........n.

              =  0             otherwise

the resulting probabilities corresponding to values of x is called Binomial Distribution. 


I. The Mean and Variance of Binomial Distribution  

i)  Mean of binomial distribution is np

ii) Variance of binomial distribution id npq

 now the Mean of the Binomial Distribution is = np = 4 x 0.6666 = 2.6664

 for calculating variance we first find the value of q, we know that  p + q = 1

therefore q = 1 - q

q = 1 - 0.6666  = 0.3334 

then the Variance of binomial distribution id npq = 4 x 0.6666 x 0.3334  = 0.8889

therefore The Mean  is 2.6664 and Variance is 0.8889 

II. P(X=2)

for finding this value we is binomial distribution we know  P(X= x) = nCx  px qn-x   

therefore in that distribution putting X=2 and the given values are n = 4, p = 0.6666 and q= 0.3334 

P ( X = 2 ) =  4C2  0.66662 0.33344-2  

P ( X = 2 ) =  {4!}/{( 2! (4 - 2)! )} x 0.4444 x 0.1111

P ( X = 2 ) = 6 x  0.04937

P ( X = 2 ) = 0.2962


III. P (X < 4) 

for finding this value we is binomial distribution we know  P(X= x) = nCx  px qn-x   

P (X < 4)  = P ( X ≤ 3) = P ( X = 1 ) +P ( X = 2) +P (X = 3)

P (X < 4) =   4C1  0.66661 0.33344-1  + 4C2  0.66662 0.33344-2  + 4C3  0.66663 0.33344-3  

P (X < 4) = 0.09881 + 0.2962 + 0.3950

P (X < 4) =  0.790

 The Mean and Variance of Binomial Distribution are 2.6664 and 0.8889 respectively, 

P ( X = 2 ) = 0.2962 and P (X < 4) =  0.790

Problems on Probability and Probability Distribution.

Example 1. if the mean and standard deviation of binomial random variable are 16 and 2 respectively. find the values of n and p.

Solution: Here given that the values of mean and standard deviation of binomial variable as

mean = 16  and S.D. = 2

we know the mean and S.D. of binomial distribution is mean = np and S.D. = √npq 

therefore mean = np =16 

and S.D. =√npq = 2 

for finding values of n and p  we firstly find variance = (S.D.)^2 = npq= (2)^2 = 4

now we take ratio of mean and variance as variance / mean = npq/np =4/16

we get q = 4/16 = 1/4 = 0.25

and we know that q = 1-p or p = 1-q = 1-0.25= 0.75

q= 0.25 and p = 0.75 putting p = 0.75 in np = 16

n x 0.75 = 16

n = 16/0.75

n = 21.3333

n = 21      because the number of observation is not expressed in fractions it is approximately 21

therefore n = 21 and p = 0.75

 Example 2. P(Ā)=0.4 and P(Bc) =0.7 find P (A U B ) if i) A & B are exclusive event and ii) A & B are independent even.

Solution: here given that P(Ā)=0.4 and P(Bc) =0.7  form this we finding P(A) and P(B)  

P(A) = 1- P(Ā) =  1-0.4 =0.6

and P(B) = 1- P(Bc)= 1-0.7 = 0.3

i) A & B are exclusive event  then P (A ∩ B ) = 0 

P (A U B ) =  P(A) +P(B) = 0.6 +0.3= 0.9

and ii) A & B are independent even. then P (A ∩ B ) = P(A) X P(B)

Therefore P (A U B ) = P(A)+P(B) - P (A ∩ B )

 P (A U B )  = P(A) + P(B) - {P(A) X P(B) }

                        = 0.6+0.3-{0.6 X 0.3)

                        = 0.9 - 0.18 

                 P (A U B )  = 0.72        

where  P(Bc) is complement event  of event B and P(Ā)  is complement event of event A

 if A & B are exclusive event then P(A U B) = 0.9 and  A & B are independent even then

 P(A U B) =0.72


Example 3. P(Ā)=0.5 and P(Bc) =0.6 find P (A U B ) if i) A & B are exclusive event and 

 ii) A & B are independent even.

Solution: here given that P(Ā)=0.4 and P(Bc) =0.7  form this we finding P(A) and P(B)  

P(A) = 1- P(Ā) =  1-0.5 =0.5

and P(B) = 1- P(Bc)= 1-0.6 = 0.4

i) A & B are exclusive event  then P (A ∩ B ) = 0 

P (A U B ) =  P(A) +P(B) = 0.5 + 0.4 = 0.9

and ii) A & B are independent even. then P (A ∩ B ) = P(A) X P(B)

Therefore P (A U B ) = P(A)+P(B) - P (A ∩ B )

 P (A U B )  = P(A) + P(B) - {P(A) X P(B) }

                        = 0.5 + 0.4 -{ 0.5 X 0.4)

                        = 0.9 - 0.20

                 P (A U B )  = 0.7       

where  P(Bc) is complement event  of event B and P(Ā)  is complement event of event A

 if A & B are exclusive event then P(A U B) = 0.9 and  A & B are independent even then

 P(A U B) =0.7

Example 4 . In Binomial distribution n = 10 and p = 0.5 then find it's mean and variance.

Solution : Given n = 10 and p = 0.5 
 we have the mean of binomial distribution is  np 
Therefore the mean = np = 10 x 0.5  = 5
and Variance of binomial distribution is  =  npq 
Therefore the variance = npq = 10x 0.5 x 0.5 = 2.5
the mean and variance of the binomial distribution is 5 and 2.5 respectively.



Comments

Post a Comment

Popular posts from this blog

B. Com. -I Statistics Practical No. 1 Classification, tabulation and frequency distribution –I: Qualitative data.

  Shree GaneshA B. Com. Part – I: Semester – I OE–I    Semester – I (BASIC STATISTICS PRACTICAL-I) Practical: 60 Hrs. Marks: 50 (Credits: 02) Course Outcomes: After completion of this practical course, the student will be able to: i) apply sampling techniques in real life. ii) perform classification and tabulation of primary data. iii) represent the data by means of simple diagrams and graphs. iv) summarize data by computing measures of central tendency.   LIST OF PRACTICALS: 1. Classification, tabulation and frequency distribution –I: Qualitative data. 2. Classification, tabulation and frequency distribution –II : Quantitative data. 3. Diagrammatic representation of data by using Pie Diagram and Bar Diagrams. 4. Graphical representation of data by using Histogram, Frequency Polygon, Frequency Curve and     Locating Modal Value. 5. Graphical representation of data by using Ogive Curves and Locating Quartile Values....

Index Number

 Index Number      Introduction  We seen in measures of central tendency the data can be reduced to a single figure by calculating an average and two series can be compared by their averages. But the data are homogeneous then the average is meaningful. (Data is homogeneous means data in same type). If the two series of the price of commodity for two years. It is clear that we cannot compare the cost of living for two years by using simple average of the price of the commodities. For that type of problem we need type of average is called Index number. Index number firstly defined or developed to study the effect of price change on the cost of living. But now days the theory of index number is extended to the field of wholesale price, industrial production, agricultural production etc. Index number is like barometers to measure the change in change in economics activities.   An index may be defined as a " specialized  average designed to measure the...

Non- Parametric Test: Run Test

Non- Parametric Test  A Non-Parametric tests is a one of the part of Statistical tests that non-parametric test does not assume any particular distribution for analyzing the variable. unlike the parametric test are based on the assumption like normality or other specific distribution  of the variable. Non-parametric test is based on the rank, order, signs, or other non-numerical data. we know both test parametric and non-parametric, but when use particular test? answer is that if the assumption of parametric test are violated such as data is not normally distributed or sample size is small. then we use Non-parametric test they can used to analyse categorical data  or ordinal data and data are obtained form in field like psychology, sociology and biology. For the analysis use the  some non-parametric test that are Wilcoxon signed-ranked test, mann-whiteny U test, sign test, Run test, Kruskal-wallis test. but the non-parametric test have lower statistical power than ...

Statistical Inference

 Statistical Inference: The Power of Statistical Inference in Data Analysis Statistical Inference: Drawing Meaningful Conclusions from Data In a data analysis, The statistical inference is a  a powerful tool for drawing meaningful conclusions from a sample of data  and making inferences about a larger population. It enables us to make confident predictions, understand relationships, and uncover valuable insights that can inform decision-making and shape various fields of study. At its core, statistical inference involves using statistical methods to analyse sample data and extend the findings to the broader population. This approach is necessary because it is often impractical or impossible to collect data from every individual or element of interest. Instead, we carefully select a representative sample and employ statistical techniques to infer information about the larger population. The first aspect of statistical inference is estimation. Estimation allows us to estima...

B. Com. I OE - II Practical No. 3 Scatter Plot.

  Practical No. 3   Scatter Plot.  Correlation In previous blog we discussed about the measure central tendency and Dispersion to use to study the variable. the correlation is a statistical concept that allows us to measure and understand the relationship between two or more variables. it provided a valuable information about that variables. e.g. price and demand of commodity, income and expenditure of family, height and weight of group of persons. their we use the relation   between this two variables. in above examples we see the one variable increases other variable is also changes in same or opposite direction.  definition: Correlation is statistical tool which study the relationship between two or more variables. for analysis of correlation various method and techniques are used.  example: i. Demand and supply of product                               ...

Median test

 Non- Parametric test Median test Median test is also a Non-Parametric test and it is alternative to Parametric T test. The median test is used when we are interested to check the two independent sample have same median or not. It is useful when data is discrete or continuous and if data is in small size.  Assumptions:  I) the variable under study is ordinal scale II) the variable is random and Independent. The stepwise procedure for computation of median test for two independent sample : Step I :- firstly we define the hypothesis Null Hypothesis is the two independent sample have same median.  Against Alternative Hypothesis is the two independent sample have different median.  Step II :- In this step we combine two sample data. And calculating the median of combined data. Step III :- after that for testing hypothesis we constructing the (2x2) contingency table. For that table we divide the sample into two parts as number of observation above and below to the ...

Non-Parametric Test 1.Sign test

Types of Non-Parametric Test : the non-parametric test is Broadly divided into three categories as: 1. One-sample Test  2.Two-sample test 3. K-sample test In this blog we discuss the one-sample tests,  there are four types of one-sample non-parametric test as follows:  1. Run Test. 2.Sign Test 3. Wilcoxon Sing-Rank test 4. Kolmogorov-smirnov test. first we discuss the sign test. 1.Sign Test The sign test id one of the very simplest non-parametric test. the test is based on signs (i.e. plus and minus signs) hence is called sign test. the non-parametric test are alternative to the parametric one, therefore the sign test is alternative to parametric T-test, is used to test the median of population instead of mean of population. (i.e. Parametric T-test is used to testing mean of population and sign test is used to testing the median of the population they are alternative to each other when the data follows assumption of T-test we use parametric t-test, but i some situation th...

Statistical Inference Practical: Point Estimation by Method of Moment

 

Statistical Inference I ( Theory of Estimation) : Unbiased it's properties and examples

 📚Statistical Inference I Notes The theory of  estimation invented by Prof. R. A. Fisher in a series of fundamental papers in around 1930. Statistical inference is a process of drawing conclusions about a population based on the information gathered from a sample. It involves using statistical techniques to analyse data, estimate parameters, test hypotheses, and quantify uncertainty. In essence, it allows us to make inferences about a larger group (i.e. population) based on the characteristics observed in a smaller subset (i.e. sample) of that group. Notation of parameter: Let x be a random variable having distribution function F or f is a population distribution. the constant of  distribution function of F is known as Parameter. In general the parameter is denoted as any Greek Letters as θ.   now we see the some basic terms :  i. Population : in a statistics, The group of individual under study is called Population. the population is may be a group of obj...

B. Com. I OE - II Practical No. 4 Karl Pearson's correlation coefficient

  Practical no. 4 Karl Pearson's correlation coefficient  Correlation In previous blog we discussed about the measure central tendency and Dispersion to use to study the variable. the correlation is a statistical concept that allows us to measure and understand the relationship between two or more variables. it provided a valuable information about that variables. e.g. price and demand of commodity, income and expenditure of family, height and weight of group of persons. their we use the relation   between this two variables. in above examples we see the one variable increases other variable is also changes in same or opposite direction.  definition: Correlation is statistical tool which study the relationship between two or more variables. for analysis of correlation various method and techniques are used.  example: i. Demand and supply of product                           ...